The magnitude of the acceleration of a particle which moves along the curve x=2 sin ¡3t,y=2 cos ¡3t and z=8t at any time t>0 is
1
Expert's answer
2019-11-04T09:21:22-0500

"a=\sqrt{(\ddot{x})^2+(\ddot{y})^2+(\ddot{z})^2}\\
\ddot{z} = 0\\
\dot{x}=6a\cos(3at)\\
\ddot{x} = -18a^2\sin(3at)\\
\dot{y}=-6a\sin(3at)\\
\ddot{y}=-18a^2\cos(3at)\\
a=\sqrt{18^2 \sin^2(3at)+18^2\cos^2(3at)}=18\sqrt{2}"